### v1

@@ 3 System architecture
Our approach comprises to main components: a tamper-proof black box-like device, and a module responsible for generating natural language explanations for the data stored by the former. The first component offers services for building, storing, and verifying integrity proofs through smart contract specifications. The explainability component provides natural language explanations by implementing Retrieval-Augmented Generation (RAG), thereby enhancing the obtained results using the data recorded by the first component as an external source to enrich the context in the answer formulation process. Both components are integrated through the existence of two main asynchronous tasks focused on cleaning and processing the raw data generated by the black box-like device to increase the effectiveness of the explainability engine.
Figure 1 depicts a generic architecture of our proposal, highlighting in blue color the functionalities related to accountability, and in green those related to explainability. Below we provide a deeper description of the approaches designed and followed in order to develop and integrate the above mentioned components.
@@ 3.1 Black Box Recorder
Our anti-tampering approach for this accountability proposal comprises two key components: a chain of integrity proofs coupled with the smart contract specification. This section delves into their design and justifies how both align with the constraints of autonomous agents.
@@ 3.1.1 Chain of integrity proofs
By chaining every hash with the previous one, we optimize the validation of stored messages due the last linked digest should correspond to the final proof saved in the blockchain. Additionally, a service has been developed to generate the chain of hashes using the recorded data, ensuring partial validation of the integrity proofs without modifying the stored messages or adding non-essential information to the transactions saved in the blockchain.
@@ 3.1.2 Smart contract
Calling a smart contract function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain. In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm (ECDSA), effectively identifying the signer.
The defined smart contract also determines its ownership, implementing access control to ensure that only the contract owner can add or update message proofs. These features guarantee authentication and non-repudiation.
Moreover, our smart contract includes a function to verify the existence of a message proof on the blockchain. This fact allows external parties to independently confirm whether a specific message hash has been stored in the contract and when it was stored. This functionality contributes to the system’s transparency, enabling users or auditors to check the validity of message proofs without relying solely on the contract owner, thereby promoting public accountability.
@@ 3.2 Explainability Component
Our solution involves use of RAG to retrieve data from outside the language model (non-parametric) and augment the prompts by adding the information recorded from our accountability component in context. Before this integration, the information undergoes processing by a dedicated module responsible for acquiring and processing the raw data generated by the autonomous agent, serving as contextual input in the question-answering system.
Given the volume of messages, their varied nature, their length, the presence of non-meaningful records generated by the autonomous agent, and the need of establishing specific relationships between messages to enhance the context of the engine, this component extracts relevant information. Its aim is to improve the effectiveness of the explanations provided to non-expert users.
@@ 4.1 Black Box Recorder
Robot Operating System (ROS) is the most commonly used framework in robotics, consisting of a set of software libraries and tools for building robot applications. However, challenges in meeting the requirements of real-time embedded systems, and the need to improve its multi-platform capabilities led to a major upgrade to ROS 2, whose main change lies in using the Data Distribution Service (DDS) (Maruyama et al., 2016).
@@
Its main functionalities include planning, control, localization, visualization and other additional features. Furthermore, it uses behavior trees to create customized navigation behavior via orchestrating many independent modular servers, which implies high flexibility in the navigation task and the specification of complex robot behaviors (Macenski et al., 2020). Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the behavior tree (BT) over a ROS 2 interface, such as an action server or service. The use of BTs to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information answer questions (Han et al., 2021).
Previous authors’ work describes the steps involved in the accountable information recording process (Fernández-Becerra et al., 2023). ROS 2 nodes publish data that other nodes consume through ROS 2 topics. Such data include perception information from sensors or commands to actuators from decision-making nodes. Navigation reproducibility of autonomous robots in ROS 2 involves storing essential data, such as their poses, maps, odometry, velocity, and laser scan readings. To this aim, we have used Rosbag files, a binary file format designed to capture and store data generated during the execution of a ROS system. These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding system behavior. The set of topics included in the Rosbag files generated are detailed in Table 1.
@@
To ensure the immutability of generated Rosbag files, Ethereum has been selected as the blockchain framework. Ethereum, an opensource public blockchain, introduces the cryptocurrency Ether and a programming platform known as Smart Contract. Employing Proof of Stake (PoS) as a consensus protocol, Ethereum offers advantages in energy efficiency, scalability, security, decentralization, accessibility, and economic efficiency. The high-level algorithms 1 and 2 depict the main steps in recording and proof verification services.
@@
Our approach addresses challenges such as the drop in robot performance during the Rosbag recording process and costs associated with storing blockchain data. We accomplish this by saving chained hashes from messages at different intervals. These intervals can be established based on the relevance of messages from a topic, regarding their required immutability features. Furthermore, message proofs are stored in real-time on the Ethereum blockchain, eliminating the need to modify the content of the Rosbag file. Detection of any alteration to the messages within this file is facilitated by a service that generates the hash chain from the file at intervals set for each topic during the recording process. To verify the existence of any of these hashes in the blockchain, the blockchain logic has been extended by developing a Smart Contract (SC) using the Solidity language. Solidity is an open-source programming language that facilitates the creation of decentralized applications, contracts, protocols, and agreements. This contract provides functionalities for storing and verifying proof information for hashes stored on the Ethereum blockchain by associating each hash value with the number of block where it is stored. Verification of transaction ownership is also conducted during the storage of these proofs.
@@ 4.2 Log Curation Component
To filter relevant information from the Rosbag file messages, a raw data processing node has been implemented. Given the size, variety, and intricate nature of messages contained in the Rosbag file, the inclusion of this raw information might be challenging for non-expert users. Therefore, it is necessary to include a component that processes this data, eliminating non-essential content in the natural language explanation generation process. By analyzing messages from the recorded Rosbag topics, this service can identify key circumstances in the navigation process, such as the appearance of an obstacle or a change in the planned trajectory. As a result, through this component, ROS 2 developers are relieved from having to call the logging API to generate explanations. This functionality streamlines the development process, allowing developers to focus on core functionalities without need for managing log calls. This approach reduces debugging efforts and boosts efficient resource usage, scalability, and adaptability in the development lifecycle, contributing to a more productive natural language explanation generation process. An example of how raw ROS messages are transformed through this component is included in Figure 2.
@@
Rosbag Raw Messages action_msgs.msg.GoalStatusArray(status_list=action_msgs.msg.GoalStatus(goal_info= action_msgs.msg.GoalInfo( goal_id=unique_identifier_msgs.msg.UUID(uuid=array([190, 81, 48, 208, 248, 147, 78, 82, 188, 116, 202, 152, 203, 225, 149, 158], dtype=uint8)), stamp=builtin_interfaces.msg.Time(sec=27, nanosec=933000000)), status=2)]) nav2_msgs.msg.BehaviorTreeStatusChange(timestamp=builtin_interfaces.msg.Time(sec=1701354686, nanosec=130104968), node_name=’ComputePathToPose’, previous_status=’IDLE’, current_status=’RUNNING’) Processed Data 1701354686 Navigation to the goal number 1 has started. 1701354686 Navigation to the goal number 1 is in progress. 1701354686 Nav2 Behavior Tree node NavigateRecovery that recovers from unexpected situations, is running.
@@ 4.3 Explainability Component
The output of the above described node serves as an external source in our implementation of the RAG approach. This development is based on the use and extension of localGPT  [^1] https://github.com/PromtEngineer/localGPT, an open-source question-answering system designed to interact with documents conversationally while safeguarding the privacy of the underlying information. Running locally, it ensures data privacy and secure interactions with the interpreted and recorded data, facilitating the generation of natural language explanations. Its functionality involves loading a set of selected documents, breaking them into smaller chunks, augmenting search capabilities with LangChain tools, and generating local embeddings. The outcomes are stored in a local vector database using Chroma vector storage. Once this information is saved, and given a user input, relevant splits are retrieved from storage via a similarity search, which finds the appropriate piece of information from the documents. The retrieved data is used by the selected LLM when producing an answer. Figure 3 depicts the main components in the RAG architecture used.
Figure 3: Retrieval Augmented Generation arquitecture.
@@ 4.4 Software and hardware artifacts
Our solution has been deployed in a Linux server running Ubuntu 22.04 as Operating System (OS). The server is equipped with 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
@@
The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotics simulations. To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services (AWS) [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world. Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik (Guzmán et al., 2017). This combination of tools and platforms enabled us to assess the impact and effectiveness of our approach.
The interaction with Ethereum has been developed through web3.py library (version 6.6.1). Ganache (version 2.7.1) has been used as the Ethereum network simulator.
RAG solution holds Langchain (version 0.325), ChromaDB (version 0.4.6). We used the open source model Llama 2 7B GGUF (Touvron et al., 2023) as the LLM responsible for generating natural language explanations for end users. This LLM integrates RLHF and shows versatility in handling general-purpose tasks, ranging from answering questions, text generation, and language translation to more demanding tasks such as reasoning or coding. Preliminary tests of Llama 2, compared to other LLMs such as Mistral (Jiang et al., 2023) or Zephyr (Tunstall et al., 2023), led us to choose this model for evaluating our data, due to the promising results obtained in these assessments conducted during the last quarter of 2023, when this work was in its development phase. As text embedding model we have chosen Instructor-xl (Su et al., 2022), a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, fact particularly important in the processes performed for any RAG solution.
@@ 5 Evaluation
To evaluate our proposal, we conducted an experimental study comparing the performance of an accountability system built in accordance with Section 4. This study covered several approaches for storing integrity proofs across various scenarios. Additionally, we explored and adapted the capabilities of RAG solutions and LLMs in question-answering systems, using as source the data gathered by our black box-like component, through a comprehensive set of questions. This work aimed to asses the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge (ERL Consumer) (ERL, 2023).
@@
Our experiments are structured around three distinct assessment scenarios in line with the navigation functionality previously described. In the initial scenario, the robot must navigate to three waypoints, encountering no obstacles along the way. The second scenario presents the robot with the challenge of adjusting its pre-planned path to reach the second goal. This adjustment becomes necessary due to the appearance of an obstacle obstructing the original planned route, requiring a deviation to achieve the goal. The third scenario introduces a new obstacle that blocks the only door needed to reach the first goal. Consequently, this obstacle prompts the cancellation of the first waypoint objective. The second and third goals are achieved with the same restrictions as detailed in the second scenario.

### v2

@@ 3 System Architecture
Our approach comprises two main components: a tamper-proof black box-like device, and a module responsible for generating natural language explanations for the data stored by the former. The first component offers services for building, storing, and verifying integrity proofs through SC specifications. The explainability component provides natural language explanations by implementing RAG, thereby enhancing the obtained results using the data recorded by the first component as an external source to enrich the context in the answer formulation process. Both components are integrated through the existence of two main asynchronous tasks focused on cleaning and processing the raw data generated by the black box-like device to increase the effectiveness of the explainability engine.
Figure 1 depicts a generic architecture of our proposal, highlighting in blue color the functionalities related to accountability, and in green those related to explainability. Below we provide a deeper description of the approaches designed and followed in order to develop and integrate the above-mentioned components.
@@ 3.1 Black Box Recorder
Our anti-tampering approach for this accountability proposal comprises two key components: a chain of integrity proofs coupled with the SC specification. This section delves into their design and justifies how both align with the constraints of autonomous agents.
@@ 3.1.1 Chain of Integrity Proofs
By chaining every hash with the previous one, we optimize the validation of stored messages because the last linked digest should correspond to the final proof saved in the blockchain. Additionally, a service has been developed to generate the chain of hashes using the recorded data, ensuring partial validation of the integrity proofs without modifying the stored messages or adding non-essential information to the transactions saved in the blockchain.
@@ 3.1.2 Smart Contract
Calling an SC function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain. In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm (ECDSA), effectively identifying the signer.
The defined SC also determines its ownership, implementing access control to ensure that only the contract owner can add or update message proofs. These features guarantee authentication and non-repudiation.
Moreover, our SC includes a function to verify the existence of a message proof on the blockchain. This fact allows external parties to independently confirm whether a specific message hash has been stored in the contract and when it was stored. This functionality contributes to the system’s transparency, enabling users or auditors to check the validity of message proofs without relying solely on the contract owner, thereby promoting public accountability.
@@ 3.2 Explainability Component
Our solution involves the use of RAG to retrieve data from outside the language model (non-parametric) and augment the prompts by adding the information recorded from our accountability component in context. Before this integration, the information undergoes processing by a dedicated module responsible for acquiring and processing the raw data generated by the autonomous agent, serving as contextual input in the question-answering system.
Given the volume of messages, their varied nature, their length, the presence of non-meaningful records generated by the autonomous agent, and the need to establish specific relationships between messages to enhance the context of the engine, the log curation component extracts relevant information. This functionality serves to enhance the effectiveness of explanations provided to non-expert users. A more comprehensive description of this component is provided in Section 4.
@@ 4.1 Black Box Recorder
ROS is the most commonly used framework in robotics, consisting of a set of software libraries and tools for building robot applications. However, challenges in meeting the requirements of real-time embedded systems, and the need to improve its multi-platform capabilities led to a major upgrade to ROS 2, whose main change lies in using the Data Distribution Service (DDS) [65].
@@
Its main functionalities include planning, control, localization, visualization and other additional features. Furthermore, it uses behavior trees to create customized navigation behavior via orchestrating many independent modular servers, which implies high flexibility in the navigation task and the specification of complex robot behaviors [66]. Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the behavior tree over a ROS 2 interface, such as an action server or service. The use of behavior trees to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information to answer questions [67].
ROS 2 nodes publish data that other nodes consume through ROS 2 topics. Such data include perception information from sensors or commands to actuators from decision-making nodes. Navigation reproducibility of autonomous robots in ROS 2 involves storing essential data, such as their poses, maps, odometry, velocity, and laser scan readings. To this aim, we have used Rosbag files, a binary file format designed to capture and store data generated during the execution of a ROS system. These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding system behavior. The set of topics included in the Rosbag files generated are detailed in Table 1.
@@
To ensure the immutability of generated Rosbag files, Ethereum has been selected as the blockchain framework. Ethereum, an open-source public blockchain, introduces the cryptocurrency Ether and self-executing programs (SCs) for code execution on the blockchain. Ethereum’s open-source nature ensures decentralization and accessibility, allowing participation without needing approval, a feature crucial for maintaining the integrity of recorded data. Additionally, Ethereum’s use of Proof of Stake (PoS) as a consensus protocol enhances energy efficiency and throughput compared to traditional protocols such as Proof of Work (PoW). In addition, Ethereum’s scalability, security and economic efficiency make it a very secure option for ensuring data immutability. The high-level algorithms 1 and 2 depict the main steps in recording and proof verification services.
@@
Our approach addresses challenges such as the drop in robot performance during the Rosbag recording process and costs associated with storing blockchain data. We accomplish this by saving chained hashes from messages at specific intervals, ensuring the integrity of data while preserving system performance. These intervals define how often an integrity proof is stored and can be established based on the relevance of messages from a topic, regarding their required immutability features. Furthermore, message proofs are stored in real-time on the Ethereum blockchain, eliminating the need to modify the content of the Rosbag file. Detection of any alteration to the messages within this file is facilitated by a service that generates the hash chain from the file at intervals set for each topic during the recording process. To verify the existence of any of these hashes in the blockchain, the blockchain logic has been extended by developing an SC using the Solidity language. Solidity is an open-source programming language that facilitates the creation of decentralized applications, contracts, protocols, and agreements. This contract provides functionalities for storing and verifying proof information for hashes stored on the Ethereum blockchain by associating each hash value with the block number where it is stored. Verification of transaction ownership is also conducted during the storage of these proofs.
@@ 4.2 Log Curation Component
To filter relevant information from the Rosbag file messages, a raw data processing node has been implemented. Given the size, variety, and intricate nature of messages contained in the Rosbag file, the inclusion of this raw information might be challenging for non-expert users. Therefore, it is necessary to include a component that processes this data, eliminating non-essential content in the natural language explanation generation process. By analyzing messages from the recorded Rosbag topics, this service can identify key circumstances in the navigation process, such as the appearance of an obstacle or a change in the planned trajectory.
The status of the navigation process is checked through the study of the ROS 2 action server NavigateToPose, used for commanding a robot to move to a specified pose. By default, Nav2 guarantees to find the shortest path under all conditions. The /plan topic includes the set of poses computed to reach a specific goal pose. If this initial path changes due to an external circumstance, the Euclidean distance between all the poses included in the pre-computed path will increase. This fact, along with the values obtained through the /scan topic, allows us to determine if the appearance of an obstacle has caused the route modification. These reasonings have been successfully checked in previous authors’ work [68] with the aim of building explanations for autonomous agents based on the use of ROS 2 topics.
Additionally, the present approach uses the information provided by Nav2 behavior trees due to their potential in designing, visualizing, and understanding the behavior of autonomous robots. Other information such as the present position of the robot and its velocity is also processed through this component. A general overview of its behavior is described through the algorithm 3. As a result, through this component, ROS 2 developers are relieved from having to call the logging API to generate explanations. This functionality streamlines the development process, allowing developers to focus on core functionalities without the need for managing log calls. This approach reduces debugging efforts and boosts efficient resource usage, scalability, and adaptability in the development lifecycle, contributing to a more productive natural language explanation generation process. An example of how raw ROS messages are transformed through this component is included in Figure 2.
Algorithm 3 Log Curation Process
1:$rosbag$, $btNodeDesc$
2:$curationLog$
3:for $message$ in $rosbag$ do
4:     $msgTopic\leftarrow$getMsgTopic($message$)
5:     $LogMsg\leftarrow message.data$
6:     // Check a change in the planned path.
7:     if $msgTopic$ == ’/plan’  then
8:         $distance\leftarrow$ getEuclideanDistance($message.poses$)
9:         if $distance>previousDistance$ DISTANCE_THRESHOLD then
10:              $logMsg\leftarrow$ “Planned path has changed when trying to achieve goal pose number $nGoal$"
11:              $previousDistance\leftarrow distance$
12:              if min($last\_scan\_msg.ranges$) $<$ OBSTACLE_DISTANCE_THRESHOLD then
13:                  $logMsg\mathrel{+}=$“ because there was an obstacle."
14:              end if
15:         end if
16:     end if
17:     // Check the navigation status.
18:     if $msgTopic$ == ’/navigate_to_pose/_action/status’  then
19:         $logMsg\leftarrow$ “No navigation is running."
20:         $statusList\leftarrow$ /navigate_to_pose/_action/status
21:         for $status$ in $statusList$ do
22:              $currentGoal\leftarrow statusList.id$
23:              if $currentGoal$ != $previousGoal$ and $currentGoal$ not in $finishedGoals$ then
24:                  $logMsg\leftarrow$ “Navigation to a new goal has started."
25:              else
26:                  $logMsg\leftarrow$ “Navigation to the goal "
27:                  if $status$ = EXECUTING then
28:                       $logMsg\mathrel{+}=$ “is in progress.”
29:                  else if $status$ = SUCCEEDED then
30:                       $logMsg\mathrel{+}=$ “has succeeded."
31:                  else if $status$ = CANCELED then
32:                       $logMsg\mathrel{+}=$ “was cancelled."
33:                  else if $status$ = ABORTED then
34:                       $logMsg\mathrel{+}=$ “has been aborted."
35:                  end if
36:                  if $status$ = SUCCEEDED or $status$ = CANCELED or $status$ = ABORTED  then
37:                       $finishedGoals\cup=currentGoal$
38:                  end if
39:              end if
40:              // Check the executed Behavior Tree nodes and their status.
41:              if $msgTopic$ == ’/behavior_tree_log’  then
42:                  $logMsg\leftarrow$“Node " $message.NodeName$ “performing "$btNodeDesc$“is in status " $message.nodeStatus$
43:              end if
44:              // Gets robot’s position and orientation.
45:              if $msgTopic$ == ’/amcl_pose’  then
46:                  $logMsg\leftarrow$ “Position: " $message.x,message.y$ “Orientation: " $message.z,message.w$
47:              end if
48:              if $msgTopic$ == ’/cmd_vel’  then
49:                  $logMsg\leftarrow$ “Linear velocity: " $message.x,message.y$ “Angular velocity: " $message.z$
50:              end if
51:         end for
52:     end if
53:     $logMsg\leftarrow message.timestamp$ + $logMsg$
54:     $curationLog\leftarrow curationLog$ + $logMsg$
55:end for
56:return $curationLog$
@@
{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Rosbag Raw Messages action_msgs.msg.GoalStatusArray(status_list=action_msgs.msg.GoalStatus(goal_info= action_msgs.msg.GoalInfo( goal_id=unique_identifier_msgs.msg.UUID(uuid=array([190, 81, 48, 208, 248, 147, 78, 82, 188, 116, 202, 152, 203, 225, 149, 158], dtype=uint8)), stamp=builtin_interfaces.msg.Time(sec=27, nanosec=933000000)), status=2)]) nav2_msgs.msg.BehaviorTreeStatusChange(timestamp=builtin_interfaces.msg.Time(sec=1701354686, nanosec=130104968), node_name=’ComputePathToPose’, previous_status=’IDLE’, current_status=’RUNNING’){mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Processed Data 1701354686 Navigation to the goal number 1 has started. 1701354686 Navigation to the goal number 1 is in progress. 1701354686 Nav2 Behavior Tree node NavigateRecovery that recovers from unexpected situations, is running.
@@ 4.3 Explainability Component
The output of the previously described node serves as an external source in our implementation of the RAG approach. This development is based on the use and extension of localGPT  [^1] https://github.com/PromtEngineer/localGPT, an open-source question-answering system designed to interact with documents conversationally while safeguarding the privacy of the underlying information. Running locally, it ensures data privacy and secure interactions with the interpreted and recorded data, facilitating the generation of natural language explanations. Its functionality involves loading a set of selected documents, breaking them into smaller chunks, augmenting search capabilities with LangChain tools, and generating local embeddings. The outcomes are stored in a local vector database using Chroma vector storage. Once this information is saved, and given a user input, relevant splits are retrieved from storage via a similarity search, which finds the appropriate piece of information from the documents. The retrieved data is used by the selected LLM when producing an answer. Figure 3 depicts the main components in the RAG architecture used.
Figure 3: Retrieval Augmented Generation architecture.
@@ 4.4 Software and Hardware Artifacts
Our solution has been deployed in a Linux server running Ubuntu 22.04 as the Operating System (OS). The server is equipped with a 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
@@
The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotics simulations. To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services (AWS) [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world. Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik [69]. This combination of tools and platforms enabled us to assess the impact and effectiveness of our approach.
The interaction with Ethereum has been developed through the web3.py library (version 6.6.1). Ganache (version 2.7.1) has been used as the Ethereum network simulator.
RAG solution holds Langchain (version 0.325), and ChromaDB (version 0.4.6). We used the open-source model Llama 2 7B GGUF [70] as the LLM responsible for generating natural language explanations for end users. This LLM integrates Reinforcement Learning from Human Feedback (RLHF) and shows versatility in handling general-purpose tasks, ranging from answering questions, text generation, and language translation to more demanding tasks such as reasoning or coding. Preliminary tests of Llama 2, compared to other LLMs such as Mistral [71] or Zephyr [72], led us to choose this model for evaluating our data, due to the promising results obtained in these assessments conducted during the last quarter of 2023, when this work was in its development phase. As a text embedding model we have chosen Instructor-xl [73], a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, a fact particularly important in the processes performed for any RAG solution.
@@ 5 Evaluation
To evaluate our proposal, we conducted an experimental study comparing the performance of an accountability system built in accordance with Section 4. This study covered several approaches for storing integrity proofs across various scenarios. Additionally, we explored and adapted the capabilities of RAG solutions and LLMs in question-answering systems, using as a source the data gathered by our black box-like component, through a comprehensive set of questions. This work aimed to assess the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge (ERL Consumer) [74].
@@
Our experiments are structured around three distinct assessment scenarios in line with the navigation functionality previously described. In the initial scenario, the robot must navigate to three waypoints, encountering no obstacles along the way. The second scenario presents the robot with the challenge of adjusting its previously planned path to reach the second goal. This adjustment becomes necessary due to the appearance of an obstacle obstructing the original planned route, requiring a deviation to achieve the goal. The third scenario introduces a new obstacle that blocks the only door needed to reach the first goal. Consequently, this obstacle prompts the cancellation of the first waypoint objective. The second and third goals are achieved with the same restrictions as detailed in the second scenario.